Kowaleski Corporation makes a product with the following standard costs: Standard Quantity or Hours Standard Price or Rate Direct materials 4.9 grams $ 7.00 per gram Direct labor 0.6 hours $ 14.00 per hour Variable overhead 0.6 hours $ 4.00 per hour In June the company produced 4,200 units using 21,830 grams of the direct material and 2,580 direct labor-hours. During the month the company purchased 24,100 grams of the direct material at a price of $6.80 per gram. The actual direct labor rate was $14.60 per hour and the actual variable overhead rate was $3.90 per hour. The materials price variance is computed when materials are purchased. Variable overhead is applied on the basis of direct labor-hours. Required: Compute the following variances for raw materials, direct labor, and variable overhead, assuming that the price variance for materials is recognized at point of purchase
Answer #1

1) Material price variance = (7-6.8)*24100 = 4820 F

Material quantity variance = (4200*4.9-21830)*7 = 8750 U

2) Labor rate variance = (14-14.60)*2580 = 1548 U

Labor efficiency variance = (4200*.6-2580)*14 = 840 U

3) Variable overhead rate variance = (4-3.90)*2580 = 258 F

Variable overhead efficiency variance = (4200*.6-2580)*4 = 240 U